Blue Hamster Manufacturing Inc. is a small firm, and several of its managers are worried about how soon the firm will be able to recover its initial investment from Project Alphas expected future cash flows. To answer this question, Blue Hamsters CFO has asked that you compute the projects payback period using the following expected net cash flows and assuming that the cash flows are received evenly throughout each year. Complete the following table and compute the projects conventional payback period. For full credit, complete the entire table Year 0 Year 1 Year 2 Year 3 Expected cash flow Cumulative cash flow 5,500,000 $2,200,000 $4,675,000 $1,925,000 Conventional payback period The conventional payback period ignores the time value of money, and this concerns Blue Hamsters CFO. He has now asked you to compute Alphas discounted payback period, assuming the company has a 8% cost of capital Complete the following table and perform any necessary calculations. Round the discounted cash flow values to the nearest whole dollar, and the discounted payback period to the nearest two decimal places. For full credit, complete the entire table. Year 0 Year 1 Year 2 Year 3 Cash flow Discounted cash flow Cumulative discounted cash flow 5,500,000 $2,200,000 $4,675,000 $1,925,000 Discounted payback periodWhich version of a projects payback period should the CFO use when evaluating Project Alpha, given its theoretical superiority? O The discounted payback period O The regular payback period One theoretical disadvantage of both payback methods-compared to the net present value method-is that they fail to consider the value of the cash flows beyond the point in time equal to the payback period How much value does the discounted payback period method fail to recognize due to this theoretical deficiency? $3,565,164 O $2,073,223 o $5,536,186 o $1,528,127
Answer #1
1) Year 0 Year 1 Year 2 Year 3
Expected cash flow -5500000 2200000 4675000 1925000
Cumulative cash flow -5500000 -3300000 1375000 3300000
Conventional payback period = 1+3300000/4675000 = 1.71 Years
Year 0 Year 1 Year 2 Year 3
2) Expected cash flow -5500000 2200000 4675000 1925000
Discounted cash flow -5500000 2037037 4008059 1528127
Cumulative discounted cash flow -5500000 -3462963 545096 2073223
Discounted payback period = 1+3462963/4008059 = 1.86 Years
3) Which version…………………………………………superiority?
*The discounted payback period.
4) Value not recognized = $2073223 (NPV)